OpenStudy (anonymous):

What is the value of the dicriminant -7x^2+2x-4=0

OpenStudy (anonymous):

how many real solutions if any?

OpenStudy (anonymous):

\[-7x^2 + 2x -4 = 0 \]

OpenStudy (anonymous):

The discriminant is given by: \[b^2 - 4ac > 0\]

OpenStudy (anonymous):

\[ax^2 + bx + c = 0\]

OpenStudy (anonymous):

Can you match up a, b and c with the coefficients of your equation?

OpenStudy (anonymous):

how many real solutions?

OpenStudy (anonymous):

i or 2 or 2 imaginary

OpenStudy (anonymous):

2

OpenStudy (anonymous):

You will retrieve 2 distinct real roots

OpenStudy (lgbasallote):

@QRAwarrior discriminant is just \(b^2 - 4ac\) ^_^ the > 0 thingies are used to find how many roots the equation has..the question is asking for the value of the discriminant

OpenStudy (anonymous):

Oops

OpenStudy (anonymous):

No but he asked how many roots

OpenStudy (anonymous):

As well

OpenStudy (lgbasallote):

oh..i just looked at the main question lol..carry on

OpenStudy (anonymous):

@lgbasallote Would you mind helping me with my question after? I think I gave the open poster sufficient amounts of help

OpenStudy (anonymous):

can u give me the final answer?

OpenStudy (anonymous):

I'm terrible at math I think someone else can really help you i'm so so sorry =-(

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

\[-7x^2 + 2x -4 = 0, then\] b = 2, a = -7, c = -4 Now, can you work out the discriminant? All you have to do is sub in the values into the following equation:

OpenStudy (anonymous):

b^2 - 4ac